OpenStudy (unklerhaukus):

\[y'={y(1-y^2) \over x(1-x^2}\]

OpenStudy (unklerhaukus):

\[{dy \over y(1-y^2)} = {dx \over x(1-x^2)}\]

OpenStudy (unklerhaukus):

\[ln(y)-{1 \over 2}ln(y^2-1)=lnx-{1 \over 2}ln(x^2-1)\]

OpenStudy (unklerhaukus):

+c

OpenStudy (unklerhaukus):

\[y^2(x^2-1)=x^2(y^2-1)k\]